Solve and check each of the following equations.
5/x+6 + 2/x^2+7x+6=3/x+1
Please help.
Thank you.

Answer by sdmmadam@yahoo.com(530)

If the problem is
5/(x+6) + 2/(x^2+7x+6)=3/(x+1)----(1)then
5/(x+6) + 2/(x+6)(x+1)= 3/(x+1) (factorising the quadratic expression)
Multiplying by (x+6)(x+1)
5(x+1)+2=3(x+6)----(2)
5x+5+2 = 3x+18
5x+7 = 3x+18
5x-3x= 18-7
2x= 11
x = 11/2
Verification:putting x = 11/2 in (2) LHS = 69/2=RHS